OpenStudy (anonymous):

Integration. How do I find the primitive \[F(x)=\int \frac{1}{(x-2)^2}\] ? Hints appreciated.

OpenStudy (raden):

int by u-sub let : u=x-2

OpenStudy (anonymous):

I would recommend you perform an u-substitution for this kind of problem: \[\Large u=x-2 \\ \Large du=dx\\ \]

terenzreignz (terenzreignz):

Before anything else, and while this may seem to be immaterial, I think it matters if you're to understand u-substitution method... It's that your integral doesn't make sense if you don't indicate by what variable you integrate. Better rewrite with the simple modification... \[F(x)=\int \frac{1}{(x-2)^2}dx\]

terenzreignz (terenzreignz):

As they suggested, you now begin the u-substitution method. You let u = x - 2 Then you get the derivative of u, with respect to x... \[\large \frac{du}{dx} = 1\] And you 'multiply both sides by dx' \[\large du = 1\cdot dx = dx\]

terenzreignz (terenzreignz):

Returning to your integral, we make the necessary substitutions, namely, replacing x-2 with u, and replacing dx by du, as we derived. \[\large \int\limits \frac{1}{u^2}du\] And now you see the importance of the differential 'du' at the integral, it tells you that now, we integrate with respect to u. Can you do it from here?

OpenStudy (anonymous):

Yes. \[\int u^{-2}du= -\frac{1}{u}+C. \text{Since u=(x-2)} ... -\frac{1}{x-2}+C\] Worked like a charm

terenzreignz (terenzreignz):

Nicely done :) I'd usually get rid of that ugly minus sign and make it \[\frac{1}{2-x}+C\] But that's fine :)